OpenStudy (anonymous):

How many solutions are there to the equation below? -46 = 3x + 7x A. Infinitely many B. 1 C. 2

OpenStudy (anonymous):

x(3+7)=-46 x*10=-46 x=-4.6 so...

OpenStudy (anonymous):

sooo

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i got x= -23/5 sooo

OpenStudy (anonymous):

@jollysailorbold same thing -23/5=-4.6

OpenStudy (anonymous):

right :)